# Spacetime foam correlation renders cosmological constant (dark energy)

### v1

@@ 1 Introduction
1.1 Wheeler spacetime foams and Planck lattice
@@
It is plausible that precisely due to the violent quantum gravitational fluctuations at the Planck length $\ell_{\rm pl}=M^{-1}_{\rm pl}$, the spacetime must exhibit the complex small-scale structure known as “foam” [1, 2, 3, 4]. This physical arena is a wormhole formed by gravitational quantum fluctuations, connecting two spacetime points with a non-trivial topology [5, 6]. The wormhole is an energetically favourable configuration that can be spontaneously nucleated in the Euclidean quantum gravity [7]. Its nucleation (creation) and absorption (annihilation) [8, 9] can be viewed as a tunnelling solution with a semi-classical amplitude proportional to the gravitational instanton action $e^{-\beta M/2}$. In this article, we consider the case that the wormhole mass $M\!\sim\!M_{\rm pl}$ and the tunnelling length $\beta\!\sim\!\ell_{\rm pl}$ between two spacetime points is of the order of the Planck unit. The physical spacetime made of fundamental “constituent” foams becomes a fluctuating foamy Planck lattice [10] endowed with a basic length $\ell_{\rm pl}$, rendering a natural ultraviolet cutoff of quantum field theories. The Planck lattice represents the ground state of Euclidean quantum gravity, studied by using the Wilson-loop area action, path-integral quantization and nonperturbative discretisation of a four-dimensional simplicial complex and Regge calculus [11, 12].
@@ 1.2 Wormhole effects on field operators at large distances
In addition to the spacetime discretisation, wormhole quantum nucleation and absorption are described by the creation and annihilation operators $\hat{\alpha}_{i}$, $\hat{\alpha}^{\dagger}_{i}$ and $[\hat{\alpha}_{i},\hat{\alpha}^{\dagger}_{i}]=\delta_{ij}$ of type-$i$ wormhole topological solution. Their interactions with field operators $O_{i}$ at distances much larger than the wormhole size at the Planck scale are [13, 14, 15, 16]
$\alpha_{i}\int\sqrt{g}d^{4}xO_{i}(x).$ (1.1)
The ground state $|\alpha_{i}\rangle$ or $\alpha_{i}$-vacuum, $(\hat{\alpha}_{i}+\hat{\alpha}^{\dagger}_{i})|\alpha_{i}\rangle=\alpha_{i}|\alpha_{i}\rangle$, is determined by a set of eigenvalues $\{\alpha_{i}\}$, representing different Universes degenerated and tunneling to each other. It is analogous to the topological $\theta$-vacuum in gauge field theories [17]. The operator $O_{i}$ subscript “$i$” indicates the type of operator associated with the type-$i$ wormhole solution. As the consequences of the interaction (1.1), the spacetime wormholes may have effects on the natural constants and fundamental interactions of effective field theories of the operators $Q_{i}$, as an example, the gravitational constant $G$, the cosmological constant $\Lambda$ and Ricci scalar $R$ of the Einstein theory. In the proposal [14, 18], the eigenvalues $\alpha_{i}$ possibility distribution and most probable value explain the vanishing of the cosmological constant, leading to an enormous spike of activity, see review [19].
Our proposal is as follows. Upon the ground state $|\alpha_{i}\rangle$ of the foamy spacetime (Planck lattice), we introduce the foamon field $\Phi_{i}$ or quantised “foamon”, a quasi-particle, to describe the collective excitations of wormhole quantum nucleation, oscillation and absorption at short distances. It is analogous to the phonon excitations of collective vibration in the nuclear lattice or the magnon excitations of the collective spin wave of electrons in a crystal lattice. This differs from the usual spacetime description $g_{\mu\nu}+h_{\mu\nu}$ by the classical metric $g_{\mu\nu}$ and quantum fluctuation $h_{\mu\nu}$. The interactions (1.1) induce the foamon field $\Phi_{i}$ coupling to physical operators $O_{i}$ as an infrared regulator at large distances. As a result, we show that the foamon field $\Phi_{i}$ correlation length renders the characteristic scale of the effective theories for the field operator $Q_{i}$. We apply this scenario to the Ricci scalar and cosmological constant of the Einstein theory for cosmology.
@@ 2.1 Foamon fields describe collective excitations of wormhole fluctuations
where $\hat{\alpha}_{i}^{\dagger}(x)$ and $\hat{\alpha}_{i}(x)$ are type “$i$” wormhole creating and annihilating operators at the point $x=(\tau,{\bf x})$, following the equal-time commutator
$[\hat{\alpha}_{i}(\tau,{\bf x}^{\prime}),\hat{\alpha}^{\dagger}_{j}(\tau,{\bf x})]=(2\pi)^{4}\delta_{ij}\delta^{3}({\bf x}^{\prime}-{\bf x}).$ (2.2)
@@
In this article, we consider the effective Lagrangian of the foamon field as  [^1] Gauge symmetries and covariant invariance preserved, and $g_{\mu\nu}$ is not explicit written.
@@
and the effective potential $V(\Phi_{i})$ represents foamon fields $\Phi_{i}$ self-interactions due to the complex dynamics of wormholes at short distances. There are possibilities for the interactions between different types “$i$” of wormholes. The potential $V(\Phi_{i})$ bounded from below, and its absolute minimum $V^{\rm min}$ is the ground state $\Phi^{\rm min}_{i}$ of the lowest energy in theory, thus, the true and stable ground state. It is located at: (i) zero field $\Phi^{\rm min}_{i}=0$, (ii) finite field $\Phi^{\rm min}_{i}\not=0$; (iii) more than two degenerate minima. We do not wander in the swamp land of possible potentials [20], but instead look at relevant and irrelevant operators in fixed points and scaling invariant domains with known interacting dynamics. Consider the simplest and $\Phi_{i}\Leftrightarrow-\Phi_{i}$ invariant potential,
@@
[…]
@@ 2.2 Foamon fields effective action and energy density
The foamon field “vacuum-to-vacuum” transition amplitude is
@@
[…]
@@ 2.3 Foamon fields correlation function and length
[…]
Here we use the translation invariance at the macroscopic length scale, which is much larger than the wormhole size, the Planck lattice spacing. In the space of the foamon field eigenstates, the diagonalized two-point function (2.10) is
@@
The fluctuating foamon fields $\Phi_{i}$ are associated with the ground state $\alpha_{i}$-vacuum. Their correlation functions (2.10) and (2.11) contain the information about the mixings (tunnelling) of different $\alpha_{i}$ vacuum states and correlations in different spacetime points.
@@ 3.1 Foamon fields interact with field operators at large distances
where the sector “$i$” is fixed, the coupling $y^{i=j}=1$ is for the same sector $i=j$ and mixing $y=y^{i\not=j}<1$ is for different sectors $i\not=j$.
@@ 3.2 Field operator at large distance acts as an infrared regulator
The interaction (3.1) can be incorporated as a generating functional
@@
Here, introducing an infrared scale $bM_{\rm pl}<M_{\rm pl}$ and a scaling parameter $b<1$, we formally split the functional modes integration $\int^{M_{\rm pl}}_{{}_{0}}=\int^{bM_{\rm pl}}_{{}_{0}}\int^{M_{\rm pl}}_{bM_{\rm pl}}$ of the foamon field $\Phi_{i}$. The infrared scale $bM_{\rm pl}<M_{\rm pl}$ represents the characteristic scale of the low-energy field operator $O_{i}$ at large distances. The scaling parameter $b_{i}$ depends on the field operator $O_{i}$ sector, and we omit its subscript ``$i^{\prime\prime}$ to simplify the notation.
@@ 3.4 Wilson approach to foamon field renormalisation
Henceforth, we discuss only one type of the foamon field and suppress the subscript “$i$” to simplify the notations. Using the Wilson renormalization approach to integrate high-energy modes $\hat{\Phi}$ (3.5) in the generating function (3.7), we can formally write
@@
indicating short distances $x^{\prime}$ (large momenta $k^{\prime}$) scale to long distances $x$ (small momenta $k$). In the coordinate space, the Wilson renormalisation approach at each step “$n$” is to obtain the mean “coarse” field $\bar{\phi}^{n+1}_{\rm coarse}=\frac{1}{V_{b}}\sum\phi^{n}_{\rm fine}$ and interactions by averaging (smearing) all “fine” fields $\phi^{n}_{\rm fine}$ and interactions in small coordinate cells $[(b^{n+1},b^{n})M_{\rm pl}]^{-4}$) over the volume $V_{b}=1/(bM_{\rm pl})^{4}$.

### v2

@@ 1 Introduction
1.1 Infrared catastrophe of Euclidean quantum gravity
The laws and constants of nature, as seen by an ordinary observer inside a macroscopic universe, relate to a background of baby universes or wormhole operators which act on a wave-function space of universes [1, 2, 3, 4, 5, 6, 7, 8, 9]. Wormholes are gravitational fluctuations of topology changes and two-point connections of the spacetime, as schematically illustrated in Fig. 1. The Coleman formulation of this idea for the cosmological constant issue [7] is based on the Euclidean path integral approach of the effective actions,
$I_{E}$ $=$ $\int\sqrt{g}d^{4}x_{1}\sqrt{g}d^{4}x_{2}O^{\dagger}_{i}(x_{1})C_{ij}(x_{1},x_{2})O_{j}(x_{2}),$ (1.1)
$\sim$ $\alpha_{i}^{\dagger}C^{-1}_{ij}\alpha_{j}+\alpha_{i}\int\sqrt{g}d^{4}xO_{i}(x),$ (1.2)
$\Rightarrow$ $\int\sqrt{g}d^{4}x\Big{[}\rho_{{}_{\Lambda}}-\frac{1}{16\pi G}R+\gamma R^{2}+\cdot\cdot\cdot\Big{]}.$ (1.3)
The two-point connection functional $C_{ij}(x_{1},x_{2})$ in the nonlocal action (1.1) represents the background wormhole dynamics, and its eigenvalues and eigenvectors relate to the auxiliary parameters $\{\alpha_{i}\}$ of the universes. The path integration (1.2) over $\{\alpha_{i}\}$ values yields the action (1.1). The various types “ii” of local scalar operators $O_{i}(x)$ represent physical observables of the universes, relating to the cosmological constant $\rho_{{}_{\Lambda}}=\Lambda/(8\pi G)$, gravitational coupling $R/G$, high-order $\gamma R^{2}$ terms of the Ricci scalar $R$, and other local field operators in the local effective action (1.3).
Figure 1: The left figure illustrates an Euclidean spacetime wormhole solution: an $R^{3}$ transits to an $R^{3}$ plus an $S^{3}$ baby universe, which is subsequently absorbed, becoming again an $R^{3}$. The right figure illustrates many wormholes in the spacetime $x$. The time in left and $t$ in right indicate the Euclidean time, i.e., the tunnelling length $\beta\sim 1/M$, and $M$ is the wormhole mass scale. We adopt these Figures from Figures 1 and 2 of the Reference [10].
The formulation [7] assumed the following background wormhole dynamics, as schematically illustrated in Fig. 2. (i) The Euclidean path integral amplitudes (probability) of different $\{\alpha_{i}\}$-parametrised universes (1.2) are dominated by the large-scale universe, connected by background wormholes $C_{ij}(x_{1},x_{2})$. (ii) The amplitude for a large-scale spherical universe is of order $\exp\{1/(G^{2}\rho_{{}_{\Lambda}})\}$ for the case of small cosmological constant $\Lambda$, that has been a priori introduced in the effective action (1.3). (iii) After integrating all wormholes connecting any pair of points $x_{1}$ and $x_{2}$, the probability of different $\alpha_{i}$-universes is given by [11]
$\sim e^{-C^{-1}_{ij}\alpha_{i}\alpha_{j}}\exp\Big{\{}\exp\frac{3}{8(\Lambda+\alpha_{1})(G+\alpha_{2})}+(\gamma+\alpha_{3})+\cdot\cdot\cdot\Big{\}},$ (1.4)
where natural constants $G$, $\Lambda$ and $\gamma,\cdot\cdot\cdot$ are functions in the $\{\alpha_{i}\}$ parameter space. Our Universe (1.3) with the observed Newton constant $G$, small cosmological constant $\Lambda$ and negligible coefficients $\gamma,\cdot\cdot\cdot$ corresponds to the particular $\{\alpha_{i}\}$ values for which the probability (1.4) is maximal.
In such a formulation, however, the amplitude runs into the instability of the infrared catastrophe, due to the absence of physically sensible infrared regulators for the background wormhole dynamics $C_{ij}$ and the unboundedness of the Euclidean action (1.3) of large-scale universes. The arbitrarily enormous numbers of macroscopic wormholes [11] and large-scale spherical universes [12] can be nucleated all the way up to the infrared scale $G^{2}\rho_{{}_{\Lambda}}\rightarrow 0$. Therefore, the amplitude runs away, and the parameters $\alpha_{i}$ are unbounded. The interpretations using the probability functional (1.4) are inconsistent. Thus, many alternative approaches studying this issue have been advocated, and they can be found in the review article [10].
Differing from the Coleman formulation, we will discuss the novel scenario of background wormhole dynamics with the physical ultraviolet and infrared regulators. As a result, we obtain a posterior the Euclidean actions (1.1) and (1.3) at long distances.
Figure 2: The Coleman scenario for background wormhole dynamics. Two (or many) large smooth geometries (spherical universes) connected by wormholes. Large universes connected by wormholes. We adopt this figure from Figure 1 of the Reference [12].
1.2 Wheeler spacetime foams and Planck lattice at short distances
As heuristically illustrated in Fig. 1, wormholes connect two spacetime points with a non-trivial topology, which gain quantitative support from a corresponding Euclidean classical solution [13]. The physical arena is of wormholes formed by gravitational (spacetime) fluctuations, connecting spacetime points with a non-trivial topology [14, 15]. Such a wormhole can be viewed as the process of the creation of an instanton (a half-wormhole [16]) and the absorption (annihilation) of an instanton, in other words, the pair creation, oscillation and annihilation of an instanton and an anti-instanton. It is analogous to the dynamics of electron and positron pairs in the QED vacuum. The gravitational instanton  [^1] For many instantons, negative Coulomb energies are present among instantons. is an energetically favourable configuration that can be spontaneously nucleated [17] in the local Euclidean quantum gravity (1.3) without the cosmological constant $\rho_{{}_{\Lambda}}$ term. Its nucleation (creation) and absorption (annihilation) [18, 19] can be viewed as a tunnelling solution with a semi-classical amplitude proportional to the gravitational instanton action $e^{-\beta M/2}$, the nucleation rate of numerous instantons is proportional to $e^{-\beta^{2}M^{2}/2}$ [17]. Here $M$ is the instanton mass (size) and $\beta$ indicates the tunnelling length between two spacetime points.
Therefore, it is a plausible conjecture that precisely due to the violent quantum gravitational fluctuations at the Planck length $\ell_{\rm pl}=M^{-1}_{\rm pl}$, numerous Planck-size wormholes (instanton and anti-instanton pairs) of mass $M\!\sim\!M_{\rm pl}$ and the tunnelling length $\beta\!\sim\!\ell_{\rm pl}$ are nucleated without exponential suppression $e^{-\beta^{2}M^{2}/2}\sim\mathcal{O}(1)$. In such a small length scale, Euclidean action well describes gravitational quantum fluctuations. As the nucleating processes are both energetically and entropically favourable, the number of Planck-size wormholes should be as large as permitted by spacetime geometry and topology, and the wormhole number density is about $\sim M^{3}_{\rm pl}$, namely the close packing density. The right of Figure 1 illustrates only a dilute density of wormholes. Such a configuration is the ground state of the spacetime that must exhibit the complex small-scale structure known as “foam” [20, 21, 22, 23].
@@ 1.2 Wheeler spacetime foams and Planck lattice at short distances
Thus, the physical spacetime made of fundamental “constituent” foams becomes a fluctuating foamy Planck lattice [24] endowed with a basic length $\ell_{\rm pl}$, rendering a natural ultraviolet cutoff of quantum field theories. The Planck lattice represents the ground state of Euclidean quantum gravity (1.3) without cosmological constant $\Lambda$. We studied it using the Wilson-loop area action, path-integral quantisation, and nonperturbative discretisation of a four-dimensional simplicial complex and Regge calculus [25, 26].
Such a speculative scenario describes the background wormholes of the spacetime. Its detailed quantum dynamics at short distances are unknown, except for providing a natural ultraviolet cutoff. However, we will discuss how long-range background wormhole dynamics couple to physical observables at long distances, which act as natural infrared regulators in the Wilson renormalisation approach.
@@ 1.3 Wormhole effects on field operators at large distances
In addition to the spacetime discretisation, instanton quantum nucleation and absorption are described by the creation and annihilation operators $\hat{\alpha}_{i}$, $\hat{\alpha}^{\dagger}_{i}$ and $\hat{\alpha}_{i},\hat{\alpha}^{\dagger}_{i}]=\delta_{ij}$ of type-$i$ wormhole topological solution. Their interactions with field operators $O_{i}$ at distances much larger than the wormhole size at the Planck scale are [8, 7, 27, 28]
$\alpha_{i}\int\sqrt{g}d^{4}xO_{i}(x),$ (1.5)
in the action (1.2). The ground state $|\alpha_{i}\rangle$ or $\alpha_{i}$-vacuum, $(\hat{\alpha}_{i}+\hat{\alpha}^{\dagger}_{i})|\alpha_{i}\rangle=\alpha_{i}|\alpha_{i}\rangle$, is determined by a set of eigenvalues $\{\alpha_{i}\}$, representing different Universes degenerated and tunneling to each other. The operators $\alpha_{i}$ and $\alpha^{\dagger}_{i}$ associate with the amplitudes $e^{-\beta M/2}$ of creation and annihilation [29], and we will adopt the approximation $e^{-\ell_{\rm pl}M_{\rm pl}/2}\sim\mathcal{O}(1)$ for Planck-size wormholes. It is analogous to the topological $\theta$-vacuum in gauge field theories [30]. The field operator $O_{i}$ subscript “ii” indicates the type of operator associated with the type-$i$ wormhole solution. As the consequences of the interaction (1.5), the spacetime wormholes may have effects on the natural constants and fundamental interactions of effective field theories of the field operators $Q_{i}$, as an example, the gravitational constant $G$, the cosmological constant $\Lambda$ and Ricci scalar $R$ of the Einstein theory. In the proposal [7, 31], the eigenvalues $\alpha_{i}$ possibility distribution and most probable value explain the vanishing of the cosmological constant, leading to an enormous spike of activity, see review [10].
Our proposal is as follows. Upon the ground state $|\alpha_{i}\rangle$ of the foamy spacetime (Planck lattice), we introduce the foamon field $\Phi_{i}$ or quantised “foamon”, a quasi-particle, to describe the collective excitations of wormhole quantum nucleation, oscillation and absorption at short distances. This foamon field represents the long-range dynamics of background wormholes. It is analogous to the phonon excitations of collective vibration in the nuclear lattice or the magnon excitations of the collective spin wave of electrons in a crystal lattice. This differs from the usual spacetime description $g_{\mu\nu}+h_{\mu\nu}$ by the classical metric $g_{\mu\nu}$ and quantum fluctuation $h_{\mu\nu}$. The interactions (1.5) induce the foamon field $\Phi_{i}$ coupling to physical operators $O_{i}$ as an infrared regulator at large distances. As a result, we show that the foamon field $\Phi_{i}$ correlation length renders the characteristic scale of the effective theories for the field operator $Q_{i}$. We apply this scenario to the Ricci scalar and cosmological constant of the Einstein theory for cosmology. As a result, we show how the long-range correlation length of background wormholes gives rise to the origin of the cosmological constant observed in the macroscopic universe.
@@ 2.1 Foamon fields describe collective excitations of wormhole fluctuations
where $\hat{\alpha}_{i}^{\dagger}(x)$ and $\hat{\alpha}_{i}(x)$ are type “ii” wormhole creating and annihilating operators at the point $x=(\tau,{\bf x})$, following the equal-time commutator
$\hat{\alpha}_{i}(\tau,{\bf x}^{\prime}),\hat{\alpha}^{\dagger}_{j}(\tau,{\bf x})]=(2\pi)^{4}\delta_{ij}\delta^{3}({\bf x}^{\prime}-{\bf x}).$ (2.2)
@@
In this article, we consider the effective Euclidean Lagrangian of the foamon field as follow  [^2] Gauge symmetries and covariant invariance preserved, and $g_{\mu\nu}$ is not explicit written.
@@
and the effective potential $V(\Phi_{i})$ represents foamon fields $\Phi_{i}$ self-interactions due to the complex dynamics of background wormholes at short distances. There are possibilities for the interactions between different types “ii” of wormholes. The potential $V(\Phi_{i})$ bounded from below, and its absolute minimum $V^{\rm min}$ is the ground state $\Phi^{\rm min}_{i}$ of the lowest energy in theory, thus, the true and stable ground state. It is located at: (i) zero field $\Phi^{\rm min}_{i}=0$, (ii) finite field $\Phi^{\rm min}_{i}\not=0$; (iii) more than two degenerate minima. We do not wander in the swamp land of possible potentials [32], but instead look at relevant and irrelevant operators in fixed points and scaling invariant domains with known interacting dynamics. Consider the simplest and $\Phi_{i}\Leftrightarrow-\Phi_{i}$ invariant potential,
@@
which preserves both local and global symmetries. In four-dimensional spacetime, the quadratic $\Phi_{i}^{2}$ is a $d=2$ relevant operator, the quartic $\Phi_{i}^{4}$ is a $d=4$ marginal operator, and the last term $(\cdot\cdot\cdot)$ indicates high-dimensional irrelevant operators. The potential (2.4) minimises $V_{\rm min}=0$ at a zero vacuum expectation value $\langle\Phi^{2}_{i}\rangle=0$, and the system is in a symmetric (disorder) phase.
@@ 2.2 Foamon fields effective action and energy density
The foamon field “vacuum-to-vacuum” transition amplitude is described by the Euclidean path integral
@@
where $\int_{x}\equiv\int_{\Omega}d^{4}x=\int_{\Omega}d\tau d^{3}x\sqrt{g}$. We define functional integration as
$\int{\mathcal{D}}\Phi=\prod_{i}\int^{M_{\rm pl}}_{0}{\mathcal{D}}\Phi_{i}=\prod_{i}\prod_{M_{\rm pl}\geq k>0}\int d\Phi_{i}(k).$ (2.6)
@@ 2.3 Foamon fields correlation function and length
$\langle\cdot\cdot\cdot\rangle$ $=$ $W^{-1}_{E}(0)\int^{M_{\rm pl}}_{{}_{0}}{\mathcal{D}}\Phi(\cdot\cdot\cdot)e^{-\int_{x}{\mathcal{L}}(\Phi_{i},\partial_{\mu}\Phi_{i})}.$
Here, we use the translation invariance at the macroscopic length scale, which is much larger than the wormhole size, the Planck lattice spacing. In the space of the foamon field eigenstates, the diagonalized two-point function (2.10) is
@@
The fluctuating foamon fields $\Phi_{i}$ are associated with the ground state $\alpha_{i}$-vacuum. Their correlation functions (2.10) and (2.11) contain the information about the mixings (tunnelling) of different $\alpha_{i}$ vacuum states and correlations in different spacetime points. By definition, the foamon fields’ correlation functions (2.10) and (2.12) differ from the connection function $C_{ij}(x_{1},x_{2})$ (1.1) of wormholes connecting two spacetime points, as shown in Fig. 2.
@@ 3.1 Foamon fields interact with field operators at large distances
where the sector “ii” is fixed, the coupling $y^{i=j}=1$ is for the same sector $i=j$ and mixing $y=y^{i\not=j}<1$ is for different sectors $i\not=j$.
@@ 3.2 Field operator at large distance acts as an infrared regulator
The interaction (3.1) and Lagrangian (2.3) can be incorporated as a generating Euclidean functional of the
@@
Here, introducing an infrared scale $bM_{\rm pl}<M_{\rm pl}$ and a scaling parameter $b<1$, we formally split the functional modes integration $\int^{M_{\rm pl}}_{{}_{0}}=\int^{bM_{\rm pl}}_{{}_{0}}\int^{M_{\rm pl}}_{bM_{\rm pl}}$ of the foamon field $\Phi_{i}$. The infrared scale $bM_{\rm pl}<M_{\rm pl}$ represents the characteristic scale of the low-energy field operator $O_{i}$ at large distances. The scaling parameter $b_{i}$ depends on the field operator $O_{i}$ sector, and we omit its subscript $i^{\prime\prime}$ to simplify the notation.
@@ 3.4 Wilson approach to foamon field renormalisation
Henceforth, we discuss only one type of the foamon field and suppress the subscript “ii” to simplify the notations. Using the Wilson renormalization approach to integrate high-energy modes $\hat{\Phi}$ (3.5) in the generating function (3.7), we can formally write
@@
indicating short distances $x^{\prime}$ (large momenta $k^{\prime}$) scale to long distances $x$ (small momenta $k$). In the coordinate space, the Wilson renormalisation approach at each step “nn” is to obtain the mean “coarse” field $\bar{\phi}^{n+1}_{\rm coarse}=\frac{1}{V_{b}}\sum\phi^{n}_{\rm fine}$ and interactions by averaging (smearing) all “fine” fields $\phi^{n}_{\rm fine}$ and interactions in small coordinate cells $(b^{n+1},b^{n})M_{\rm pl}]^{-4}$) over the volume $V_{b}=1/(bM_{\rm pl})^{4}$.